If x + 3 = a, then 2x+ 6 =
(A) a + 3
(B) a + 6
(C) 2a
(D) 2a + 3
(E) 2a + 6

Answer by checkley75(3666)

x+3=a & 2x+6=2(x+3)
thus 2x+6=2a or answer c)